OpenStudy (anonymous):

(-10,3) m=3/5 point slope into slope intercept how do i do this??

OpenStudy (anonymous):

point slope is y-y1=m(x-x1) (x1,y1) is the point given. (-10,3) and m is the slope. 3/5 just plug those in and simplify to get slope intercept

OpenStudy (anonymous):

yeah i after i got point slope i tried simplifying and i couldnt get past a certain part

OpenStudy (anonymous):

Show me your work so far

OpenStudy (anonymous):

im stuck on this 3/5x +y=9

OpenStudy (anonymous):

Hmm y-3=3/5(x+10) y-3=3/5x+6 y=3/5x+9

OpenStudy (anonymous):

ugh thanks ill look at what i did wrong

OpenStudy (anonymous):

Can you see or want me to do it more detailed and show each step

OpenStudy (anonymous):

i got it, i probably just maybe a minor mistake like i always do